OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 7x + 10 = 0? No real solutions Two irrational solutions Two different rational solutions Two identical rational solutions

OpenStudy (anonymous):

Is that an x^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you must use the quadratic equation, have you learned it yet?

OpenStudy (anonymous):

No not yet

OpenStudy (anonymous):

to find the values of your x's use the quadratic equation: x=(-b +or- sqrt(b^2 -4ac)) / 2a Your equation is: x2 + 7x + 10 = 0 Your A = 1 (first coefficient of x^2) Your B = 7 (from 7x) Your C = 10 Plug those into the quadratic equation up top doing it with the + and then the - and you get two answers. then figure out which one of those 4 options they are

OpenStudy (anonymous):

\[(-b+\sqrt{b^2 - 4ac})/2a\]

OpenStudy (anonymous):

\[(-b-\sqrt{b^2 - 4ac})/2a\]

OpenStudy (anonymous):

I'm kinda confused...but would it be A?

OpenStudy (anonymous):

There's definitely a solution..plug in your values into the equation..